# Artificial Intelligence for Operations Research: Revolutionizing the Operations Research Process

### v1

@@ 6.1 Enhancement for gradient-based methods
    where $G_{t}$ is a rough approximation of the diagonal elements of the Hessian matrix and the symbol *** is multiplication between a scalar and a vector.
@@
Framework
Paper
Methodology
LSTM-based
Andrychowicz et al., (2016)
First propose a basic LSTM-based framework.
Lv et al., (2017)
Use random scaling and convex regularizers to stabilize the training process.
Chen et al., 2020a
Use curriculum learning to gradually increase the maximal number of iterations.
Chen et al., 2020b
Learn a variational stopping policy to determine the stopping time.
Metz et al., (2019)
Use the weighted average of two unbiased gradient estimators.
Wichrowska et al., (2017)
Enhance the RNN structure in a hierarchical manner.
RL-based
Li and Malik, (2016)
First propose a basic RL-based framework.
Li and Malik, (2017)
Combine the coordinates with permutation invariance property into one group.
Zheng et al., (2022)
Learn an agent that outputs a symbolic gradient update formula.
@@ 6.3 Enhancement for column-generation methods
where the binary variables $y_{p}$ decide whether the column $p$ should be selected for the next RMP. More specifically, after solving problem (6.3), if $y_{p}=1$, then column $p$ is a candidate for selection, otherwise, column $p$ is excluded. The hyperparameter $\lambda$ in (16a) balances the trade-off between minimizing the [RMP] objective and controlling the size of $\mathcal{G}$. In other words, by solving Problem (6.3), we can derive a better subset $\mathcal{G}$ with two advantages. Firstly, the subsequent [RMP] objective is greatly reduced, leading to an improved solution for the master problem. Secondly, the size of the resulting subset remains manageable. Once the MILP is solved, the better subset $\mathcal{G}$ is determined by:
@@
However, solving a MILP at each iteration can be time-consuming in practice. Thus the authors suggest imitating this MILP using AI techniques. More specifically, they represent the MILP as a bipartite graph as the one introduced in Section 2.1 (seen in Figure 2). Then they train a standard GNN to imitate the optimal $\{y_{p}\}_{p\in\bar{\mathcal{G}}}$ from the MILP (6.3). This is a binary classification task. In the inference stage, the GNN will predict the probability that a generated column $p$ should be selected, i.e., $Pr(y_{p}=1)$. If this probability is greater than $0.5$, then column $p$ is considered promising and added to the [RMP] in the next iteration.
@@ 7.1.1 Variable selection
Paper
Goal
Methodology
Features
Di Liberto et al., (2016)
Learn to switch between different variable selection strategies
Cluster MILP problems into clusters using the K-means algorithm and learn an assignment of branching methods to each cluster
40 features describing different statistics of the problem
Balcan et al., (2018)
Learn to combine the scores returned by different existing variable selection strategies
Empirical risk minimization to find the optimal weight for convex combinations
Similar to algorithm configuration, it finds an optimal weight for a certain distribution of MILPs. Since the weight only applies to this distribution, it requires no instance-specific features.
Khalil et al., (2016)
Learn to mimic the strong branching strategy
Imitation learning via SVM
18 static features and 54 dynamic features
Alvarez et al., (2017)
Learn to mimic the strong branching strategy
Imitation learning via random forests
static problem features, dynamic problem features and dynamic optimization features
Gasse et al., (2019)
Learn to mimic the strong branching strategy
Imitation learning using GNN
Bipartite graph representation with 5 features for the constraint, 13 features for the variable and 1 feature for the constraint matrix
Gupta et al., (2020)
Learn to mimic the strong branching strategy
Imitation learning using GNN at the root node and SVM at remaining nodes
same features as Gasse et al., (2019) at the root node and same features as Khalil et al., (2016) at remaining nodes
Gupta et al., (2022)
Learn to mimic the strong branching strategy
Imitation learning via GNN and Lookback regularization
same features as Gasse et al., (2019)
Zarpellon et al., (2021)
Learn to mimic the SCIP’s default branching strategy
Imitation learning using deep neural network
25 features from the candidate variables and 61 features describing the state of the B&B search tree
Sun et al., (2020)
Learn a novel branching strategy
Reinforcement learning with GNN
same features as Gasse et al., (2019)
Qu et al., (2022)
Learn a novel branching strategy
Reinforcement learning with Double DQN
same features as Gasse et al., (2019)
@@ 7.1.2 Node selection
Paper
Goal
Methodology
Features
He et al., (2014)
Learn to mimic a theoretically optimal node selection strategy which knows the optimal solution in advance and only expands nodes containing the optimal solution
Imitation learning via SVM
node features, branching features and B&B tree features
Song et al., (2018)
Same as He et al., (2014)
Retrospective imitation learning
node features and tree features
Yilmaz and Yorke-Smith, (2021)
Learn to mimic the SCIP’s default node selection strategy
Use imitation learning to learn an operator that can decide which child node to expand
Same as Gasse et al., (2019)
Labassi et al., (2022)
Same as He et al., (2014)
Imitation learning via GNN
Same as Gasse et al., (2019)
@@ 7.4 Heuristics
AI-based heuristics often draw inspiration from traditional heuristics. In the following, we first discuss the heuristics inspired by category 1) and then by category 2). Within category 1), two AI-based heuristics have been proposed inspired by FP and diving respectively. Qi et al., (2021) propose a smart feasibility pump (SFP) method inspired by the traditional feasibility pump (FP) heuristic. This new method surpasses the traditional one by reducing the number of steps needed to reach the first feasible solution. FP first solves the relaxed problem (7.1). Its solution ${x}^{(0)}$ is then rounded to obtain an initial integer solution $\bar{x}^{(0)}$. However, due to rounding, this solution is usually infeasible w.r.t. constraints (17b) and (17c). FP iteratively projects the solution $\bar{x}^{(t)}$ back to the feasible region of the relaxed problem (7.1), obtaining $x^{(t+1)}$ and rounding it to $\bar{x}^{(t+1)}$. The projection wants to satisfy constraints (17b) to (17c) with the smallest change on the current solution, while the rounding focuses only on the integer constraints. This process may take a long time since the projection and rounding focus on short-term benefits. Furthermore, there is a risk of stagnation, i.e., failing to find any feasible solution. In contrast, SFP trains an RL agent such that the infeasibility will rapidly decrease in several iterations. SFP inherits the framework of FP, but the key point is that SFP allows the agent to change the current solution $x^{(t)}$ before rounding. The reward is defined as negative infeasibility. The cumulative rewards will drive the agent to balance between current-step and long-term negative infeasibility. The state includes the information of Problem (7.1) and the current solution. The action of the agent is the change $a^{(t)}$ in the current solution $x^{(t)}$. Specifically, the state transition incorporates the change and rounding, i.e., $x^{(t+1)}=[x^{(t)}+a^{(t)}]$, where []delimited-[][$\cdot$][ ⋅ ] is the rounding operation. Despite the non-differentiable nature of rounding and infeasibility with respect to the agent’s actions, the PPO algorithm (See section 2.3) can still be effectively utilized, since it is designed to handle such a challenge.
@@ 8 Conclusion
Acknowledgements
Bissan Ghaddar’s research is supported by the Natural Sciences and Engineering Research Council of Canada Discovery Grant 2017-04185 and by the David G. Burgoyne Faculty Fellowship.

### v2

@@ 6.1 Enhancement for gradient-based methods
    where $G_{t}$ is a rough approximation of the diagonal elements of the Hessian matrix and the symbol * is multiplication between a scalar and a vector.
@@
Framework Paper Methodology
LSTM-based Andrychowicz et al., (2016) First propose a basic LSTM-based framework.
Lv et al., (2017) Use random scaling and convex regularizers to stabilize the training process.
Chen et al., 2020a Use curriculum learning to gradually increase the maximal number of iterations.
Chen et al., 2020b Learn a variational stopping policy to determine the stopping time.
Metz et al., (2019) Use the weighted average of two unbiased gradient estimators.
Wichrowska et al., (2017) Enhance the RNN structure in a hierarchical manner.
RL-based Li and Malik, (2016) First propose a basic RL-based framework.
Li and Malik, (2017) Combine the coordinates with permutation invariance property into one group.
Zheng et al., (2022) Learn an agent that outputs a symbolic gradient update formula.
@@ 6.3 Enhancement for column-generation methods
where the binary variables $y_{p}$ decide whether the column $p$ should be selected for the next RMP. More specifically, after solving problem (16), if $y_{p}=1$, then column $p$ is a candidate for selection, otherwise, column $p$ is excluded. The hyperparameter $\lambda$ in (16a) balances the trade-off between minimizing the [RMP] objective and controlling the size of $\mathcal{G}$. In other words, by solving Problem (16), we can derive a better subset $\mathcal{G}$ with two advantages. Firstly, the subsequent [RMP] objective is greatly reduced, leading to an improved solution for the master problem. Secondly, the size of the resulting subset remains manageable. Once the MILP is solved, the better subset $\mathcal{G}$ is determined by:
@@
However, solving a MILP at each iteration can be time-consuming in practice. Thus the authors suggest imitating this MILP using AI techniques. More specifically, they represent the MILP as a bipartite graph as the one introduced in Section 2.1 (seen in Figure 2). Then they train a standard GNN to imitate the optimal $\{y_{p}\}_{p\in\bar{\mathcal{G}}}$ from the MILP (16). This is a binary classification task. In the inference stage, the GNN will predict the probability that a generated column $p$ should be selected, i.e., $Pr(y_{p}=1)$. If this probability is greater than $0.5$, then column $p$ is considered promising and added to the [RMP] in the next iteration.
@@ 7.1.1 Variable selection
Paper Goal Methodology Features
Di Liberto et al., (2016) Learn to switch between different variable selection strategies Cluster MILP problems into clusters using the K-means algorithm and learn an assignment of branching methods to each cluster 40 features describing different statistics of the problem
Balcan et al., (2018) Learn to combine the scores returned by different existing variable selection strategies Empirical risk minimization to find the optimal weight for convex combinations Similar to algorithm configuration, it finds an optimal weight for a certain distribution of MILPs. Since the weight only applies to this distribution, it requires no instance-specific features.
Khalil et al., (2016) Learn to mimic the strong branching strategy Imitation learning via SVM 18 static features and 54 dynamic features
Alvarez et al., (2017) Learn to mimic the strong branching strategy Imitation learning via random forests static problem features, dynamic problem features and dynamic optimization features
Gasse et al., (2019) Learn to mimic the strong branching strategy Imitation learning using GNN Bipartite graph representation with 5 features for the constraint, 13 features for the variable and 1 feature for the constraint matrix
Gupta et al., (2020) Learn to mimic the strong branching strategy Imitation learning using GNN at the root node and SVM at remaining nodes same features as Gasse et al., (2019) at the root node and same features as Khalil et al., (2016) at remaining nodes
Gupta et al., (2022) Learn to mimic the strong branching strategy Imitation learning via GNN and Lookback regularization same features as Gasse et al., (2019)
Zarpellon et al., (2021) Learn to mimic the SCIP’s default branching strategy Imitation learning using deep neural network 25 features from the candidate variables and 61 features describing the state of the B&B search tree
Sun et al., (2020) Learn a novel branching strategy Reinforcement learning with GNN same features as Gasse et al., (2019)
Qu et al., (2022) Learn a novel branching strategy Reinforcement learning with Double DQN same features as Gasse et al., (2019)
@@ 7.1.2 Node selection
Paper Goal Methodology Features
He et al., (2014) Learn to mimic a theoretically optimal node selection strategy which knows the optimal solution in advance and only expands nodes containing the optimal solution Imitation learning via SVM node features, branching features and B&B tree features
Song et al., (2018) Same as He et al., (2014) Retrospective imitation learning node features and tree features
Yilmaz and Yorke-Smith, (2021) Learn to mimic the SCIP’s default node selection strategy Use imitation learning to learn an operator that can decide which child node to expand Same as Gasse et al., (2019)
Labassi et al., (2022) Same as He et al., (2014) Imitation learning via GNN Same as Gasse et al., (2019)
@@ 7.4 Heuristics
AI-based heuristics often draw inspiration from traditional heuristics. In the following, we first discuss the heuristics inspired by category 1) and then by category 2). Within category 1), two AI-based heuristics have been proposed inspired by FP and diving respectively. Qi et al., (2021) propose a smart feasibility pump (SFP) method inspired by the traditional feasibility pump (FP) heuristic. This new method surpasses the traditional one by reducing the number of steps needed to reach the first feasible solution. FP first solves the relaxed problem (18). Its solution ${x}^{(0)}$ is then rounded to obtain an initial integer solution $\bar{x}^{(0)}$. However, due to rounding, this solution is usually infeasible w.r.t. constraints (17b) and (17c). FP iteratively projects the solution $\bar{x}^{(t)}$ back to the feasible region of the relaxed problem (18), obtaining $x^{(t+1)}$ and rounding it to $\bar{x}^{(t+1)}$. The projection wants to satisfy constraints (17b) to (17c) with the smallest change on the current solution, while the rounding focuses only on the integer constraints. This process may take a long time since the projection and rounding focus on short-term benefits. Furthermore, there is a risk of stagnation, i.e., failing to find any feasible solution. In contrast, SFP trains an RL agent such that the infeasibility will rapidly decrease in several iterations. SFP inherits the framework of FP, but the key point is that SFP allows the agent to change the current solution $x^{(t)}$ before rounding. The reward is defined as negative infeasibility. The cumulative rewards will drive the agent to balance between current-step and long-term negative infeasibility. The state includes the information of Problem (17) and the current solution. The action of the agent is the change $a^{(t)}$ in the current solution $x^{(t)}$. Specifically, the state transition incorporates the change and rounding, i.e., $x^{(t+1)}=[x^{(t)}+a^{(t)}]$, where []delimited-[][$\cdot$][ ⋅ ] is the rounding operation. Despite the non-differentiable nature of rounding and infeasibility with respect to the agent’s actions, the PPO algorithm (See section 2.3) can still be effectively utilized, since it is designed to handle such a challenge.